# Anticipating Future Object Compositions without Forgetting

### v1

@@ 1 Introduction
Although humans have never seen a blue apple, they can easily picture it. This is due to the inherent human ability to generalize to novel concepts by combining the known entity "apple" with the color "blue." However, do computer vision models possess this capability? This question has motivated the development of Compositional Zero-Shot Learning (CZSL) [5, 21, 22, 26]. In CZSL, the goal is to recognize unseen object-attribute combinations, referred to as compositions, based on the compositions seen during training. For this, models should understand the attributes and objects that compose these compositions to generalize to all possible compositions.
@@
Unfortunately, these models tend to be biased towards object categories rather than attributes, which makes it suffer from feature misalignment when used directly for attribute recognition [3]. Fine-tuning VLMs can solve this but often leads to catastrophic forgetting of prior knowledge [28], thereby compromising their generalization ability. To address this, we explore how to fine-tune VLMs to perform well in CZSL without forgetting any prior knowledge. Nayak et al. [15] introduced Compositional Soft Prompting (CSP), which combats catastrophic forgetting by adding auxiliary tokens for all words in a given dataset and training only these tokens. CSP improves model performance in CZSL for image classification. We incorporate CSP in Grounding DINO, to leverage it for object detection.
We consider CSP as a baseline and improve it for CZSL by introducing Compositional Anticipation (CA). CA consists of two components: Compositional Smoothing and Compositional Independence. Compositional Smoothing anticipates novel object-attribute compositions by assigning soft labels when predictions are partially correct, e.g. the object is correct but the attribute is different. This approach deviates from conventional Label Smoothing [20], which assigns soft labels to all classes. Compositional Independence disentangles objects from attributes through Separation and Decorrelation. Separation introduces a separation loss to maximize the distinction between object and attribute classes by applying intra-class separation within objects and attributes and an inter-class separation between objects and attributes. Decorrelation minimizes the correlation between objects and attributes to reduce dependency between the two.
@@ 3.6 Compositional Anticipation
Our method, which we refer to as Compositional Anticipation (CA), consists of both Compositional Smoothing and Compositional Independence. Figure 1 shows how we implement CA in Grounding DINO [10].
@@ 4.4 Improving Incrementally
Figure 2 shows the effects of Compositional Prompt Tuning on the model’s predictions. Figure 2a shows that before incremental learning, 39% of all instances of "green cylinder" are misclassified as "green cube", and all instances of "brown sphere" are misclassified as "yellow sphere". Figure 2b demonstrates that after applying Compositional Prompt Tuning, "green cube," "green cylinder," "yellow sphere," and "brown sphere" are classified correctly on all instances.
@@ 5 Conclusion
Given these findings, we recommend conventional fine-tuning of Grounding DINO for applications where performance on a specific dataset is prioritized, and our method for scenarios emphasizing overall performance across datasets.
Considering that Grounding DINO excels in CZSL, likely due to the cross-modality fusion between image and text embeddings and our proposed methods involve anticipating how the embeddings should be positioned in the embedding space. Having demonstrated the benefits of our approach for CZSL, further investigation into the positioning of embeddings in the fused embedding space could potentially yield results approaching those achieved by conventional fine-tuning of Grounding DINO, but without encountering catastrophic forgetting.

### v2

@@ 1 Introduction
Although humans have never seen a blue apple, they can easily picture it. This is due to the inherent human ability to generalize to novel concepts by combining the known entity "apple" with the color "blue". However, do computer vision models possess this capability? This question has motivated the development of Compositional Zero-Shot Learning (CZSL) [5, 21, 22, 26]. In CZSL, the goal is to recognize unseen object-attribute combinations, referred to as compositions, based on the compositions seen during training. For this, models should understand the attributes and objects that compose these compositions to generalize to all possible compositions.
@@
Unfortunately, these models tend to be biased towards object categories rather than attributes, which makes it suffer from feature misalignment when used directly for attribute recognition [3]. Fine-tuning VLMs can solve this but often leads to catastrophic forgetting of prior knowledge [28], thereby compromising their generalization ability. To address this, we explore how to fine-tune VLMs to perform well in CZSL without forgetting any prior knowledge. Nayak et al. [15] introduced Compositional Soft Prompting (CSP), which combats catastrophic forgetting by adding auxiliary tokens for all words in a given dataset and training only these tokens. This approach preserves the model’s original embeddings, allowing it to retain and revert to its initial knowledge when necessary, unlike full fine-tuning, which alters the model’s parameters. CSP improves model performance in CZSL for image classification. We incorporate CSP in Grounding DINO, to leverage it for object detection.
We consider CSP as a baseline and improve it for CZSL by introducing Compositional Anticipation (CA), which recognizes that additional compositions may exist beyond those present during training. In this context, the term "anticipation" does not refer to actively predicting new compositions. Instead, it involves enhancing the model’s ability to handle potential new compositions by adjusting how it processes partially correct predictions through Compositional Smoothing and by guiding the model to disentangle attributes from objects via Compositional Independence. Compositional Smoothing prepares the model for novel compositions by assigning soft labels when predictions are partially correct, e.g., the object is correct but the attribute is different. This approach deviates from conventional Label Smoothing [20], which assigns soft labels to all classes. Compositional Independence disentangles objects from attributes through Separation and Decorrelation. Separation introduces a separation loss to maximize the distinction between object and attribute classes by applying intra-class separation within objects and attributes and an inter-class separation between objects and attributes. Decorrelation minimizes the correlation between objects and attributes to reduce dependency between the two.
@@ 3.6 Compositional Anticipation
Our method, which we refer to as Compositional Anticipation (CA), consists of both Compositional Smoothing and Compositional Independence. While CA does not actively predict unseen compositions, it prepares the model by refining how it handles potential new compositions and disentangles attributes from objects. Figure 1 shows how we implement CA in Grounding DINO [10].
@@ 4.4 Improving Incrementally
Figure 2 shows the effects of Compositional Prompt Tuning on the model’s predictions. Figure 2a shows that before incremental learning, 39% of all instances of "green cylinder" are misclassified as "green cube", and all instances of "brown sphere" are misclassified as "yellow sphere". Figure 2b demonstrates that after applying Compositional Prompt Tuning, "green cube", "green cylinder", "yellow sphere", and "brown sphere" are classified correctly on all instances.
@@ 5 Conclusion
Given these findings, we recommend conventional fine-tuning of Grounding DINO for applications where performance on a specific dataset is prioritized, and our method for scenarios emphasizing overall performance across datasets. However, we acknowledge that our experiments are limited to the CLEVR dataset, and it remains unclear how the proposed methods will perform on real-world datasets beyond this toy dataset.
Considering that Grounding DINO excels in CZSL, likely due to the cross-modality fusion between image and text embeddings and our proposed methods involve strategically guiding the positioning of embeddings in the embedding space. Having demonstrated the benefits of our approach for CZSL, further investigation into the positioning of embeddings in the fused embedding space could potentially yield results approaching those achieved by conventional fine-tuning of Grounding DINO, but without encountering catastrophic forgetting.